# Should I Share this Translation? Evaluating Quality Feedback for User Reliance on Machine Translation

### v1

@@ Acknowledgements
We would like to thank the members of the clip lab at University of Maryland for their constructive feedback and support. Specifically, we are grateful to Hiba El Oirghi, Yu Hou, HyoJung Han, Kartik Ravisankar, Nishant Balepur, Navita Goyal, and Calvin Bao for their help in testing the user interface. We especially appreciate Yimin Xiao for extensive discussions on the analysis in the paper. This work was supported in part by the Human Language Technology Center of Excellence at Johns Hopkins University, by NSF Fairness in AI Grant 2147292, and by the Institute for Trustworthy AI in Law and Society (TRAILS), which is supported by the National Science Foundation under Award No. 2229885. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of NSF or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for governmental purposes notwithstanding any copyright annotation therein.
@@
  • Carton et al. (2020) Samuel Carton, Anirudh Rathore, and Chenhao Tan. 2020. Evaluating and characterizing human rationales. arXiv preprint arXiv:2010.04736.
@@
  • Fernandes et al. (2025) Patrick Fernandes, Sweta Agrawal, Emmanouil Zaranis, André F. T. Martins, and Graham Neubig. 2025. Do llms understand your translations? evaluating paragraph-level mt with question answering. Preprint, arXiv:2504.07583.
@@
  • Ki et al. (2025) Dayeon Ki, Kevin Duh, and Marine Carpuat. 2025. Askqe: Question answering as automatic evaluation for machine translation. Preprint, arXiv:2504.11582.
@@ A.1 LLM Explanation
We show prompt used for generating explanations with LLaMA-3 70b (Grattafiori et al., 2024).
@@ B.1 Gold Annotation
We provide details of the gold annotation process used to collect gold shareability labels. We use Qualtrics [^11] https://www.qualtrics.com to design the survey and Prolific [^12] https://www.prolific.com to recruit annotators fluent in both English and Spanish. A total of 40 examples, each consisting of an English sentence and its Spanish translation, were presented in randomized order. As shown in Figure 6, annotators were asked to judge whether the Spanish translation was 1) Safe to share as-is or 2) Needs bilingual review before sharing. The survey took a median completion time of 30 minutes. We recruited 5 annotators and compensated each with 8 USD (equivalent to 16 USD/hour).
@@ B.2 Data Selection Criteria
We ensure the selected examples from the ContraTICO dataset (Ki et al., 2025) meet the following criteria before running the gold annotation: 1) Examples have a balanced distribution across three error severity levels: no error, minor errors, and critical errors in MT; 2) Examples have comparable lengths across error severity levels to minimize the influence of sentence length on participants’ confidence (Zouhar and Bojar, 2020); 3) Examples are relevant to the scenario context. We focus on the Wikivoyage [^13] https://www.wikivoyage.org/ subset of the dataset, which contains announcements and protocols related to COVID-19 (e.g., ‘‘It was declared a pandemic by the World Health Organization (WHO) on 11 March 2020.’’) instead of informal conversations (e.g., ‘‘and does this pain move from your chest?’’) or sentences with highly technical terms (e.g., ‘‘Like all coronaviruses, virions consist of single-stranded positive-sense RNA enclosed within an envelope.’’).
@@ C.1 Annotation Interface
We built a custom annotation interface, with screenshots shown in Figure 7 following the task flow: 1) Consent to Participate, 2) Pre-task survey, 3) Task instructions and compensation details, 4) Tutorial, 5) Independent decision-making task, 6) AI-assisted decision-making task, and 7) Post-task survey. Participants were required to answer all pre- and post-task survey questions. We present an interactive tutorial before the main task to ensure that participants understand the questions they will be asked for each example. We also illustrate how each type of quality feedback was presented during the AI-assisted decision-making step.
@@ C.2 Recruitment Process
Приложение D Computational & Time Efficiency
@@
LLM Explanation LLaMA-3 (70B) $\times$ 8 01:56
@@
QA Table LLaMA-3 (70B) $\times$ 8 03:48
Таблица 2: Average computational and time efficiency for four types of quality feedback. Computation is reported in GPU units (RTX A5000). Avg. Time is shown in MM:SS.
Приложение E Detailed Results
E.1 Independent vs. AI-Assisted
@@ E.2 Switch Percentage
Using this, we categorize decision transitions as shown in Table 3: 1) Over-reliance: changing from a correct to an incorrect decision after feedback (TP $\rightarrow$ FP and TN $\rightarrow$ FN); 2) Under-reliance: maintaining an incorrect decision after feedback (FP $\rightarrow$ FP and FN $\rightarrow$ FN); 3) Appropriate reliance: either maintaining a correct decision (TP $\rightarrow$ TP and TN $\rightarrow$ TN) or correcting an incorrect one after feedback (FP $\rightarrow$ TP and FN $\rightarrow$ TN).
@@
Таблица 3: Categorization of decision transitions from Independent (rows) to AI-Assisted (columns). Appropriate: Appropriate reliance; Over: Over-reliance; Under: Under-reliance.
@@
Таблица 4: Detailed switch percentage results by quality feedback type and reliance (appropriate, over-, under-reliance).
E.3 Per Shareability Label
@@ E.4 Participant Responses
Таблица 5: Detailed results for decision accuracy and CWA by quality feedback type. Ind.: Independent decision-making step; AI: AI-assisted decision-making step. Values represent mean scores with corresponding 95% confidence intervals. Best scores for each column is bold.
@@
Таблица 6: Detailed results for decision accuracy and CWA by quality feedback type and shareability label. [Uncaptioned image]: Safe to share as-is; [Uncaptioned image]: Needs bilingual review before sharing. Values represent mean scores with corresponding 95% confidence intervals. Best scores for each column is bold.
@@
Таблица 7: Participants’ responses on how they used quality feedback in their decision-making process, specifically why the feedback was perceived as confusing or unhelpful ([Uncaptioned image]) or why it was considered helpful ([Uncaptioned image]).
@@
(a) Consent to Participate
(b) Pre-task Survey
(c) Task Instructions
(d) Tutorial
(e) Independent Decision-Making
(f) AI-Assisted: Refer to caption Error Highlights
@@
(h) AI-Assisted: Refer to caption LLM Explanation
(i) AI-Assisted: Refer to caption Backtranslation
(j) AI-Assisted: Refer to caption QA Table
(k) Attention Check Question
(l) Post-task Survey

### v2

@@ Acknowledgements
We would like to thank the members of the clip lab at University of Maryland for their constructive feedback and support. Specifically, we are grateful to Hiba El Oirghi, Yu Hou, HyoJung Han, Kartik Ravisankar, Nishant Balepur, Navita Goyal, and Calvin Bao for their help in testing the user interface. We especially appreciate Yimin Xiao for extensive discussions on the statistical analysis in the paper. This work was supported in part by the Human Language Technology Center of Excellence at Johns Hopkins University, by NSF Fairness in AI Grant 2147292, and by the Institute for Trustworthy AI in Law and Society (TRAILS), which is supported by the National Science Foundation under Award No. 2229885. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of NSF or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for governmental purposes notwithstanding any copyright annotation therein.
@@
  • Bo et al. (2025) Jessica Y Bo, Sophia Wan, and Ashton Anderson. 2025. To rely or not to rely? evaluating interventions for appropriate reliance on large language models. In Proceedings of the 2025 CHI Conference on Human Factors in Computing Systems, CHI ’25, New York, NY, USA. Association for Computing Machinery.
@@
  • Carton et al. (2020) Samuel Carton, Anirudh Rathore, and Chenhao Tan. 2020. Evaluating and characterizing human rationales. In Proceedings of the 2020 Conference on Empirical Methods in Natural Language Processing (EMNLP), pages 9294–9307, Online. Association for Computational Linguistics.
@@
  • Deutsch et al. (2021) Daniel Deutsch, Tania Bedrax-Weiss, and Dan Roth. 2021. Towards question-answering as an automatic metric for evaluating the content quality of a summary. Transactions of the Association for Computational Linguistics, 9:774–789.
@@
  • Fernandes et al. (2025) Patrick Fernandes, Sweta Agrawal, Emmanouil Zaranis, Andre Martins, and Graham Neubig. 2025. Do LLMs understand your translations? evaluating paragraph-level MT with question answering. In Second Conference on Language Modeling.
@@
  • Guo et al. (2024) Ziyang Guo, Yifan Wu, Jason Hartline, and Jessica Hullman. 2024. A decision theoretic framework for measuring AI reliance. pages 221–236.
  • Guo et al. (2025a) Ziyang Guo, Yifan Wu, Jason Hartline, and Jessica Hullman. 2025a. Unexploited information value in human-AI collaboration. Preprint, arXiv:2411.10463.
  • Guo et al. (2025b) Ziyang Guo, Yifan Wu, Jason Hartline, and Jessica Hullman. 2025b. The value of information in human-AI decision-making. Preprint, arXiv:2502.06152.
@@
  • Ki et al. (2025) Dayeon Ki, Kevin Duh, and Marine Carpuat. 2025. AskQE: Question answering as automatic evaluation for machine translation. In Findings of the Association for Computational Linguistics: ACL 2025, pages 17478–17515, Vienna, Austria. Association for Computational Linguistics.
@@
  • Rajpurkar et al. (2016) Pranav Rajpurkar, Jian Zhang, Konstantin Lopyrev, and Percy Liang. 2016. SQuAD: 100,000+ questions for machine comprehension of text. In Proceedings of the 2016 Conference on Empirical Methods in Natural Language Processing, pages 2383–2392, Austin, Texas. Association for Computational Linguistics.
@@
  • Schemmer et al. (2022) Max Schemmer, Patrick Hemmer, Niklas Kühl, Carina Benz, and Gerhard Satzger. 2022. Should i follow AI-based advice? measuring appropriate reliance in human-AI decision-making. Preprint, arXiv:2204.06916.
@@
  • Srinivasan and Thomason (2025) Tejas Srinivasan and Jesse Thomason. 2025. Adjust for trust: Mitigating trust-induced inappropriate reliance on AI assistance. Preprint, arXiv:2502.13321.
@@ A.1 LLM Explanation
We show prompt used for generating explanations with LLaMA-3.3 70B (Grattafiori et al., 2024).
@@ B.1 Gold Annotation
We provide details of the gold annotation process used to collect gold shareability labels. We use Qualtrics [^12] https://www.qualtrics.com to design the survey and Prolific [^13] https://www.prolific.com to recruit annotators fluent in both English and Spanish. A total of 40 examples, each consisting of an English sentence and its Spanish translation, were presented in randomized order. As shown in Figure 6, annotators were asked to judge whether the Spanish translation was (1) Safe to share as-is or (2) Needs bilingual review before sharing. The survey took a median completion time of 30 minutes. We recruited 5 annotators and compensated each with 8 USD (equivalent to 16 USD/hour).
@@ B.2 Data Selection Criteria
We ensure the selected examples from the ContraTICO dataset (Ki et al., 2025) meet the following criteria before running the gold annotation: (1) Examples have a balanced distribution across three error severity levels: no error, minor errors, and critical errors in MT; (2) Examples have comparable lengths across error severity levels to minimize the influence of sentence length on participants’ confidence (Zouhar and Bojar, 2020); (3) Examples are relevant to the scenario context. We focus on the Wikivoyage [^14] https://www.wikivoyage.org/ subset of the dataset, which contains announcements and protocols related to COVID-19 (e.g., ‘‘It was declared a pandemic by the World Health Organization (WHO) on 11 March 2020.’’) instead of informal conversations (e.g., ‘‘and does this pain move from your chest?’’) or sentences with highly technical terms (e.g., ‘‘Like all coronaviruses, virions consist of single-stranded positive-sense RNA enclosed within an envelope.’’).
@@ C.1 Annotation Interface
We built a custom annotation interface, with screenshots shown in Figure 7 following the task flow: (1) Consent to Participate, (2) Pre-task survey, (3) Task instructions and compensation details, (4) Tutorial, (5) Independent decision-making task, (6) AI-assisted decision-making task, and (7) Post-task survey. Participants were required to answer all pre- and post-task survey questions. We present an interactive tutorial before the main task to ensure that participants understand the questions they will be asked for each example. We also illustrate how each type of quality feedback was presented during the AI-assisted decision-making step.
@@ C.2 Recruitment Process
Приложение D Inferring AI Feedback Decisions
During dataset selection, we ensure that each feedback condition includes a relatively balanced number of examples reflecting both correct and incorrect AI predictions. Since none of the feedback types provide a direct prediction of shareability judgment, we apply tailored proxy measures to approximate the intended shareability label for each feedback type, as detailed below.
Error Highlights.
For each error annotation, we take the highest error severity level h (No error, Minor, Major, or Critical) as the QE model’s decision. The decision is mapped to the two-point shareability scale as follows:
[…]
LLM Explanation.
We prompt the same model used to generate textual explanations (LLaMA-3.3 70B) to make shareability judgments using the same information and option format provided to participants in the task instructions. Exact prompt used is outlined in Appendix A.1.
Backtranslation.
Following Ki et al. (2025), we compute xCOMET-QE scores between the Spanish MT and its backtranslation. We fit a three-component Gaussian Mixture Model (GMM). [^15] https://scikit-learn.org/stable/modules/mixture.html GMM is a probabilistic clustering model that assumes data points are generated from a mixture of Gaussian distributions, assigning a probability to each point for belonging to a specific cluster. We hypothesize two clusters corresponding to different ranges of QE scores: (1) [Uncaptioned image] ‘‘Safe to share as-is’’ cluster with higher QE scores and (2) [Uncaptioned image] ‘‘Needs bilingual review before sharing’’ cluster with lower QE scores. Each of the 20 examples is assigned a probability to each cluster and is categorized into the cluster with the highest probability.
QA Table.
We follow a similar procedure to that used for backtranslation. However, instead of using QE scores, we compute AskQE scores through a two-step process: (1) measuring overlap between answers from the source $A_{\mathrm{src}}$ and from the backtranslated MT output $A_{\mathrm{bt}}$, and (2) aggregating question-answer similarities into a segment-level metric. We use F1, a standard similarity measure in QA research (Rajpurkar et al., 2016; Deutsch et al., 2021). Formally, for each example $e\in E$, we compute:
$\text{AskQE}(e)=\sum_{i=1}^{E}{\frac{\mathrm{F1}(A_{\mathrm{src}},A_{\mathrm{bt}})}{E}}$ (4)
We fit a two-component Gaussian Mixture Model (GMM) to the AskQE scores and assign each example to the cluster with the highest probability.
The selected examples were then manually reviewed by the authors to ensure that the proxy measures provided a reasonable approximation of the feedback’s implied shareability judgments. This process resulted in a relatively balanced distribution of correct and incorrect examples across feedback types: we included 9, 10, 11, and 9 correct examples for the Error Highlights, LLM Explanation, Backtranslation, and QA Table conditions, respectively (with corresponding incorrect counts of 11, 10, 9, and 10). This helped ensure that each feedback condition reflected a comparable underlying error rate.
Приложение E Computational & Time Efficiency
@@
LLM Explanation LLaMA-3.3 (70B) $\times$ 8 01:56
@@
QA Table LLaMA-3.3 (70B) $\times$ 8 03:48
Таблица 2: Average computational and time efficiency for four types of quality feedback. Computation is reported in GPU units (RTX A5000). Avg. Time is shown in $M$:$S$.
Приложение F Potential Learning Effects
While our original two-step setup (Independent → AI-Assisted) follows a widely adopted setup in prior work on AI-assisted decision making (Guo et al., 2024, 2025a, 2025b), particularly in studies focused on measuring reliance and trust in AI systems (Buçinca et al., 2021; Schemmer et al., 2022; Mehandru et al., 2023; Bo et al., 2025; Srinivasan and Thomason, 2025), we conducted an additional human study to directly address the concern of potential practice effects. In this true repetition control condition (Independent → Independent), participants completed the same task without receiving any quality feedback between steps. Specifically, we conducted (1) a pre-screening survey to recruit participants matching the original user demographics (English as first/primary/fluent language, low or no Spanish fluency, and residing in the U.S.), and (2) the main task, where each participant viewed the same 20 examples twice in succession and made shareability and confidence judgments both times without any feedback in between.
We recruited 18 participants for this control condition, all of whom self-reported English as their first language, with an average English proficiency of 5/5 and an average Spanish proficiency of 1.83/5. The task took a median of 12 minutes and 30 seconds to complete.
As shown in Table 3, both decision accuracy and confidence-weighted accuracy (CWA) in Step 2 for the true control group are significantly lower than those in all other feedback conditions (M = 0.605, S.E. = 0.036 for accuracy; M = 0.172, S.E. = 0.054 for CWA). Additionally, the control group’s performance in Step 1 did not differ significantly from that of the other condition groups. These findings suggest that the gains observed in the feedback conditions are not attributed to repeated exposure to the same example, but rather to the quality feedback provided in Step 2. Moreover, the within-group difference between Step 1 and Step 2 in the true control group showed no statistically significance, indicating that simply seeing the same example again without any feedback does not lead to better decision-making nor appropriate reliance.
In the post-survey, we asked the true control group participants to rate their experience on two dimensions: (1) Mental workload: ‘‘How much mental burden (e.g., thinking, deciding) did you experience while doing the task?’’, and (2) Helpfulness: ‘‘How helpful was seeing the same example twice in your Spanish translation quality assessment?’’ Compared to the feedback conditions, the true control group reported higher mental burden (M = 5.83) and lower perceived helpfulness (M = 2.89), suggesting that repeated exposure without any quality feedback increased cognitive load without enhancing perceived utility.
Group Condition Step 1 Acc. (↑) Step 2 Acc. (↑) Step 1 CWA (↑) Step 2 CWA (↑) Metal burden (1-7, ↓) Helpfulness (1-5, ↑)
True control - 0.613 (0.036) 0.605 (0.037) 0.182 (0.055) 0.172 (0.054) 5.83 2.89
Condition Error Highlights 0.593 (0.022) 0.647 (0.025) 0.156 (0.041) 0.239 (0.042) 4.94 3.83
LLM Explanation 0.634 (0.023) 0.693 (0.024) 0.232 (0.042) 0.353 (0.042) 4.06 4.39
Backtranslation 0.603 (0.024) 0.716 (0.025) 0.188 (0.041) 0.388 (0.043) 4.06 4.39
QA Table 0.659 (0.024) 0.765 (0.022) 0.282 (0.040) 0.488 (0.040) 4.00 4.39
Таблица 3: Comparison of the true control group and two-step groups across quality feedback types. Reported are mean values (with standard errors) for decision accuracy and CWA, as well as average ratings for two post-task survey questions: mental burden and helpfulness.
Приложение G Detailed Results
G.1 Independent vs. AI-Assisted
@@ G.2 Switch Percentage
Using this, we categorize decision transitions as shown in Table 4: (1) Over-reliance: changing from a correct to an incorrect decision after feedback (TP $\rightarrow$ FP and TN $\rightarrow$ FN); (2) Under-reliance: maintaining an incorrect decision after feedback (FP $\rightarrow$ FP and FN $\rightarrow$ FN); (3) Appropriate reliance: either maintaining a correct decision (TP $\rightarrow$ TP and TN $\rightarrow$ TN) or correcting an incorrect one after feedback (FP $\rightarrow$ TP and FN $\rightarrow$ TN).
@@
Таблица 4: Categorization of decision transitions from Independent (rows) to AI-Assisted (columns). Appropriate: Appropriate reliance; Over: Over-reliance; Under: Under-reliance.
@@
Таблица 5: Detailed switch percentage results by quality feedback type and reliance (appropriate, over-, under-reliance).
G.3 Per Shareability Label
@@ G.4 Participant Responses
Таблица 6: Detailed results for decision accuracy and CWA by quality feedback type. Ind.: Independent decision-making step; AI: AI-assisted decision-making step. Values represent mean scores with corresponding 95% confidence intervals. Best scores for each column is bold.
@@
Таблица 7: Detailed results for decision accuracy and CWA by quality feedback type and shareability label. [Uncaptioned image]: Safe to share as-is; [Uncaptioned image]: Needs bilingual review before sharing. Values represent mean scores with corresponding 95% confidence intervals. Best scores for each column is bold.
@@
Таблица 8: Participants’ responses on how they used quality feedback in their decision-making process, specifically why the feedback was perceived as confusing or unhelpful ([Uncaptioned image]) or why it was considered helpful ([Uncaptioned image]).
@@
(a) Pre-task Survey
(b) Task Instructions
(c) Tutorial
(d) Independent Decision-Making
(e) AI-Assisted: Refer to caption Error Highlights
(f) AI-Assisted: Refer to caption LLM Explanation
@@
(h) AI-Assisted: Refer to caption Backtranslation
(i) AI-Assisted: Refer to caption QA Table
(j) Attention Check Question
(k) Post-task Survey